Siri plans to retire in 4 years with 466,000 dollars in her account, which has an annual return of 4.23 percent. If she receives annual payments of X, with her first payment of X received in 4 years and her last payment of X received in 9 years, then what is X, the amount of each payment?
Answer #1

Payment of X received in year 4. So n=   4
  
Payment of X received in year 9. So n=   9
Interest rate(i)=   4.23%


Retirement fund (Present value)=   466000
today Present value is Equal to sum of Present values of Future money received  
466000 = Present value of sums received   

Present value of Future value= FV/(1+i)^n  


So 466000 = X/(1+4.23%)^4+ X/(1+4.23%)^9  
466000 = (X*0.8472840851)+ (X*0.6887559444)  
466000 = X*   1.536040029
X = 466000/ 1.536040029
X=   303377.5104
So each payment is $303377.51